OpenStudy (anonymous):

evaluate the integral: ∫sin⁡2x/√(cos^2 x+16) dx

sam (.sam.):

\[\int\limits \frac{\sin(2x)}{\cos^2x+16}dx\]

sam (.sam.):

Hmm try subbing \[u=\cos^2x+16\] \[du=-2\sin(x)\cos(x)dx\] \[-du=2\sin(x)\cos(x)dx\] From \[\int\limits \frac{\sin(2x)}{\cos^2x+16}dx\] \[=\int\limits \frac{2\sin(x)\cos(x)}{\cos^2x+16}dx\] Substituting gives \[=-\int\limits \frac{1}{u} \, du\]

OpenStudy (amriju):

there's root in the denom. i think

sam (.sam.):

Like this \[\int\limits \frac{ \sin (2 x)}{\sqrt{ \cos ^2x+16}} \, dx\]

OpenStudy (amriju):

yep..

OpenStudy (amriju):

ok...so sin 2x= 2sinx cosx..

OpenStudy (amriju):

let cos x be t...-sin xdx=dt..

OpenStudy (amriju):

then u'll find the expression reducing to t/(t^2+16)^1/2..

OpenStudy (amriju):

now 2tdt=d(t^2)..

OpenStudy (amriju):

u'll then get d(k)/(k+16)^1/2...k=t^2...

sam (.sam.):

You don't have to do that, take the whole thing from the denominator, \[u=\cos^2x+16\] \[-du=2\cos(x)\sin(x)dx\] Which gives \[\int\limits\frac{-du}{\sqrt u}\]

OpenStudy (amriju):

I didn't mention the constants to be multiplied or divided

OpenStudy (amriju):

lol yeah

sam (.sam.):

Yeah so just integrate and sub back u, done